# Search Arena: Analyzing Search-Augmented LLMs

### v1

@@ B.3 Topic Modeling
To analyze topic diversity across benchmarks, we apply the BERTopic framework [19] to the 11,764 unique English prompts in the Search Arena dataset. We generate prompt embeddings using OpenAI’s text-embedding-3-large model, perform dimensionality reduction and clustering via UMAP and HDBSCAN, and summarize each resulting cluster using GPT-4o [53]. We adapt the Arena Explorer [53] methodology for the Search Arena dataset. Figure A7 illustrates the distribution of topics derived from BERTopic clustering over the Search Arena prompts. The prominence of categories such as Technology Comparisons, Market Analysis, and Entertainment Characters, alongside a long-tail of niche domains, highlights the dataset’s breadth and its suitability for evaluating search-augmented LLMs under diverse topics.
Figure A6: Search Arena Conversation Length (Number of Turns) Distribution. Search Arena chats are multi-turn with 22.4% of conversations containing more than 1 turn.
Figure A7: Top Topic Categories in Search Arena. This figure shows the distribution of topic clusters. The most prevalent topics include Technology Comparisons (22.0%), Market Analysis (12.3%), and Entertainment Characters (10.6%). Less frequent but still diverse topics include health and shopping. The long-tail distribution reflects the breadth of real-world usage of search-augmented LLMs.
Figure A8: Search Arena Language Distribution (top 50). Search Arena prompts span 71 languages.
@@ Appendix C Describing Differences
To extract interpretable differences across two text corpora (e.g., prompts, responses), we use LLM-based dataset differencing methods [72, 14].
@@ C.1 Prompt Differences
[…]
@@ C.2 Response Differences
[…]
@@ D.1 Leaderboard Analysis
Models supported on the Search Arena platform are shown in Table A1. Number of battles across models is shown in Figure A9; pairwise battle count is shown in Figure A10.
Figure A9: Battle Count Distribution. Number of battles across Search Arena models. The distribution is not even due to (1) different sampling weights and (2) models were not added to the platforms at the same time.
Figure A10: Pairwise Battle Count Distribution. Number of battles between Search Arena models.
Average win rates across models are shown in Figure A11. Pairwise win rates are shown in Figure A12.
Figure A11: Win Rate Distribution. Average win rates of Search Arena models.
Figure A12: Pairwise Win Rates. Pairwise win rates between Search Arena models.
Consistent with Chatbot Arena’s ranking system [10], we use Bradley-Terry regression [5] to calculate model coefficients and then re-scale to match the scale of the Elo rating system. The resulting model scores are shown in Figure A13.
Figure A13: Search Arena Leaderboard. Model scores based on Elo-scaled Bradley-Terry coefficients.
Search Arena model ratings and ranks are shown in Table A3. We also compute model scores and the leaderboard for two subsets of the full dataset:
@@
Table A3: Search Arena Leaderboard. Elo-scaled Bradley-Terry ratings along with corresponding ranks (Rating (Rank)). Model ratings based on two subsets of the full dataset: (1) English vs non-English prompts, (2) Factual (Factual Lookup and Info Synthesis) vs non-Factual prompts.
@@ D.2 Response Length Analysis
Average response length distribution across Search Arena models is shown in Figure A14 (Left). Reasoning models tend to be more verbose except for sonar-reasoning. sonar-pro’s version with a higher search context generates longer responses compared to the version with a medium context. Response length distribution is uniform across OpenAI’s models.
Figure A14: (Left) Response length distribution. Reasoning models and models with high search context size tend to be more verbose. (Right) Citation count distribution. As expected, models with higher search context size cite more sources. Reasoning models cite fewer sources compared to non-reasoning variants (e.g., sonar-reasoning-pro-high vs sonar-pro-high).
We control for response length difference in the Bradley-Terry model and compute the corresponding coefficient across different subsets of the full data, broken down by intent category (see Figure A15). For all intent categories, the coefficient (effect) is statistically significant; however, as expected, the effect is smallest for Factual Lookup prompts.
Figure A15: Response Length Control across Intents. Bradley-Terry coefficients corresponding to response length across different intent categories. Length has less effect on Factual Lookup prompts.
@@ E.1 Citation Count
Citation count distribution is shown in Figure A14 (Right). As expected, models with higher search context size cite more sources (e.g., sonar-pro-high vs sonar-pro). Furthermore, reasoning models tend to cite less sources compared to non-reasoning variants (e.g., sonar-reasoning-pro-high vs sonar-pro-high, sonar-reasoning vs sonar). We hypothesize that reasoning models synthesize and filter irrelevant sources before final response generation, resulting in less cited sources in the final response (see Figure 3). Interestingly, gemini-2.5-pro-grounding cites fewer sources compared to gemini-2.5-flash-grounding, suggesting that even though both are reasoning models, gemini-2.5-pro-grounding filters out more sources from the final response.
We then control for citation count difference in the Bradley-Terry model and compute the corresponding coefficient across different subsets of the dataset broken down by intent category (see Figure A16). The effect of citation count on Guidance (e.g., debugging, problem-solving) prompts is not significant. The effect is largest on Analysis prompts.
Figure A16: Citation Count Control across Intents. Bradley-Terry coefficients corresponding to citation count across different intent categories. Citation count does not have significant effect on Guidance (e.g., debugging, problem-solving) prompts. The effect is largest for Analysis queries.
@@ E.2 Citation Sources
Figure A17: Citation Source Distribution across Model Families. Models from different providers are biased towards different types of sources: (1) Perplexity models prefer citing YouTube, social media, and community blogs, (2) OpenAI models are biased towards mainstream news outlets.
The distribution of citation domain categories is shown in Figure A17. Perplexity’s models tend to cite social media (including YouTube) and community platforms (e.g., Reddit, Quora). OpenAI’s GPT-4o-based models are more biased towards mainstream news outlets. Google’s Gemini models are in between.
We can also control for all features simultaneously (response length, number of citations, and citation sources) to compute adjusted model scores. The scores and rankings before and after applying these controls are shown in Figure A18. We observe that the model scores and rankings tend to converge after the controls are applied. This convergence is particularly evident within model families: the top three models from Perplexity show significant convergence. This suggests that much of the variation between models in the same family is accounted for by the control features. We note that features such as factuality and coverage may be correlated with the number of citations or specific citation sources; we leave this analysis to future work.
Figure A18: Model Scores Before and After Control. Model scores and rankings converge after the controls are applied.
@@ E.3 Citation Attribution Analysis
[…]
@@ Appendix F Cross-Setting Analysis
For offline evaluation of models in search settings, we use BrowseComp [61] and SimpleQA [60] [^2] Due to the high cost of running search-augmented LLMs, we evaluated each model on the same randomly sampled subset of 500 questions. The subset was sampled once and shared across all models.. For testing the general performance of the models in non-search settings, we use ArenaHard-v2 [35]. BrowseComp and SimpleQA assess factuality based on well-specified ground truths, while ArenaHard-v2 utilizes an LLM-as-a-judge framework on a set of challenging prompts filtered from Chatbot Arena’s Text Arena dataset. Although these benchmarks test different aspects of standard and search-augmented LLMs, all are partially captured by the Search Arena prompt distribution shown in Figure 1.
@@
The selected models have near-zero accuracy on the BrowseComp benchmark; this finding is not surprising, as the questions are specifically designed to evaluate and challenge deep research pipelines. The model scores for the rest of the benchmarks–Search Arena, SimpleQA, and ArenaHard-v2–are provided in Table A4. Additionally, we calculate model win rates on two subsets of the Search Arena dataset based on the annotated intent classes (see Section 2)–Search Arena (Fact+Synth) includes only Factual Lookup and Info Synthesis queries, while Search Arena (Other) contains the remaining subset. We use Kendall’s tau for comparing ranks and Pearson correlation for comparing raw scores across benchmarks. On SimpleQA, model accuracy is saturated (ranging from 89% to 93%), with minimal separability between the models and low agreement with Search Arena ($\tau=0.422$, $r=0.582$). In contrast, ArenaHard-v2 shows greater performance variance and separability, and has a higher agreement ($\tau=0.556$, $r=0.844$) with Search Arena compared to SimpleQA. However, the rankings differ; notably, the three reasoning models rank lower in ArenaHard-v2, suggesting that, while combining web search with reasoning improves performance in Search Arena, it may degrade performance on ArenaHard-v2 prompts. Additionally, when comparing the two subsets of the Search Arena with ArenaHard, Search Arena (Other) has higher agreement ($\tau=0.644$, $r=0.882$) compared to Search Arena (Fact+Synth) ($\tau=0.511$, $r=0.675$). This finding is expected as the prompt distribution in Search Arena (Other) is closer to that of ArenaHard (e.g., creative writing, problem-solving).
@@
Furthermore, we compare model performance with and without search on the SimpleQA and ArenaHard-v2 benchmarks. We used Gemini models in this case study, as search is implemented as a tool (GoogleSearch tool) and can be easily turned on and off. The performance change is shown in Figure A20. Search significantly improves model performance on SimpleQA (models’ performance saturates at around 90%). However, performance degrades on ArenaHard-v2, with non-search variants achieving higher scores.
@@
Table A4: Model scores across three benchmarks–Search Arena (win rate), SimpleQA (accuracy), and ArenaHard-v2 (win rate). On SimpleQA, models’ performance is saturated and lead to minimal separability in model scores. Search Arena and ArenaHard-v2 provide more separability, but the rankings are different.
@@
Figure A19: Model Scores Across Benchmark. (1) Models’ performance on SimpleQA is saturated with minimal separability. (2) Performance variance in ArenaHard-v2 is comparable to that of Search Arena, but the rankings are different, with reasoning models having higher performance in search settings. (3) Model scores and ordering in Search Arena (Other) is closer to that of ArenaHard compared to Search Arena (Fact+Synth).
Figure A20: SimpleQA vs ArenaHard-v2 Performance. Search and non-search performance of Gemini models on SimpleQA and ArenaHard-v2 benchmarks. Search improves performance on SimpleQA, while degrades the score on ArenaHard-v2.

### v2

@@ B.3 Topic Modeling
To analyze topic diversity across benchmarks, we apply the BERTopic framework (Grootendorst, 2022) to the 11,764 unique English prompts in the Search Arena dataset. We generate prompt embeddings using OpenAI’s text-embedding-3-large model, perform dimensionality reduction and clustering via UMAP and HDBSCAN, and summarize each resulting cluster using GPT-4o (Tang et al., 2025). We adapt the Arena Explorer (Tang et al., 2025) methodology for the Search Arena dataset. Figure B.4 illustrates the distribution of topics derived from BERTopic clustering over the Search Arena prompts. The prominence of categories such as Technology Comparisons, Market Analysis, and Entertainment Characters, alongside a long-tail of niche domains, highlights the dataset’s breadth and its suitability for evaluating search-augmented LLMs under diverse topics.
Figure B.3: Search Arena Conversation Length (Number of Turns) Distribution. Search Arena chats are multi-turn with 22.4% of conversations containing more than 1 turn.
Figure B.4: Top Topic Categories in Search Arena. This figure shows the distribution of topic clusters. The most prevalent topics include Technology Comparisons (22.0%), Market Analysis (12.3%), and Entertainment Characters (10.6%). Less frequent but still diverse topics include health and shopping. The long-tail distribution reflects the breadth of real-world usage of search-augmented LLMs.
Figure B.5: Search Arena Language Distribution (top 50). Search Arena prompts span 71 languages.
@@ Appendix C Describing Differences
To extract interpretable differences across two text corpora (e.g., prompts, responses), we use LLM-based dataset differencing methods (Zhong et al., 2022; Dunlap et al., 2024).
@@ C.1 Prompt Differences
LLM Prompt for Proposing Distinguishing Properties between Prompt Sets The following are a two separate lists of prompts that users have asked to a chatbot: <START> {text} <END> I am a machine learning researcher trying to figure out the major differences between these two groups of prompts. This is a very small portion of the data, so I want the differences to be general. Please provide a list of the top three differences (separated by bullet points "*") between the prompts in Group A and Group B. Follow the detailed instructions below: - Identify properties that are more common in Group A prompts compared to Group B prompts. - The property descriptions should be simple and detailed. Do not produce vague and generic descriptions. - Start the description of each property with "User". Do not use keywords like "group A", "group B", "more common", "less common", "frequently", "occasionally" in your response. - Do not combine multiple properties / features in a single bullet point. - An example of the desired output format: * "Property 1" * "Property 2" * "Property 3" Please order your response in terms of the most common differences between the two groups. Your response:
LLM Prompt for Reducing Proposed Properties PROPERTIES: ############################## {properties} ############################## Above is a list of properties, several of which are similar. - Please reduce this to a list of the five most common distinct properties, ordered by frequency of occurrence. - There should be no similar / overlapping properties in the final list. - Do not group multiple unrelated properties / features into a single property. - Do not use keywords, like "frequent", "rare", "common", "uncommon", etc. - Only respond with a numbered list of properties, no other text.
LLM Prompt for Ranking Properties PROMPT: <START> {text} <END> PROPERTY: <START> {hypothesis} <END> Check whether the PROMPT satisfies the PROPERTY. Respond with Yes or No. If you are unsure, respond with No. Output:
@@ C.2 Response Differences
LLM Prompt for Proposing Distinguishing Properties between Model A and Model B Responses The following is a list of Model A and Model B outputs to a set of user questions: <START> {text} <END> I am a machine learning researcher trying to figure out the major differences between the outputs of model A and model B. This is a very small portion of the data, so I want the differences to be general. Please provide a list of three top differences (separated by bullet points "*") between the outputs of model A and model B. Follow the detailed instructions below: - Identify properties that are more common in model A outputs compared to model B outputs. - The property descriptions should be simple and detailed. Do not produce vague and generic descriptions. - Do not use keywords like "model A", "model B", "more common", "less common", "frequently", "occasionally" in your response. - Do not combine multiple properties / features in a single bullet point. - An example of the desired output format: * "Property 1" * "Property 2" * "Property 3" Please order your response in terms of the most common differences between the two groups. Your response:
LLM Prompt for Ranking Properties QUESTION: <START> {question} <END> ANSWER: <START> {answer} <END> PROPERTY: <START> {hypothesis} <END> Check whether the ANSWER satisfies the PROPERTY. Respond with Yes or No. If you are unsure, respond with No. Output:
@@ D.1 Leaderboard Analysis
Models supported on the Search Arena platform are shown in Table A.1. Number of battles across models is shown in Figure D.1; pairwise battle count is shown in Figure D.2.
Figure D.1: Battle Count Distribution. Number of battles across Search Arena models. The distribution is not even due to (1) different sampling weights and (2) models were not added to the platforms at the same time.
Figure D.2: Pairwise Battle Count Distribution. Number of battles between Search Arena models.
Average win rates across models are shown in Figure D.3. Pairwise win rates are shown in Figure D.4.
Figure D.3: Win Rate Distribution. Average win rates of Search Arena models.
Figure D.4: Pairwise Win Rates. Pairwise win rates between Search Arena models.
Consistent with Chatbot Arena’s ranking system (Chiang et al., 2024), we use Bradley-Terry regression (Bradley & Terry, 1952) to calculate model coefficients and then re-scale to match the scale of the Elo rating system. The resulting model scores are shown in Figure D.5.
Figure D.5: Search Arena Leaderboard. Model scores based on Elo-scaled Bradley-Terry coefficients.
Search Arena model ratings and ranks are shown in Table D.1. We also compute model scores and the leaderboard for two subsets of the full dataset:
@@
Table D.1: Search Arena Leaderboard. Elo-scaled Bradley-Terry ratings along with corresponding ranks (Rating (Rank)). Model ratings based on two subsets of the full dataset: (1) English vs non-English prompts, (2) Factual (Factual Lookup and Info Synthesis) vs non-Factual prompts.
@@ D.2 Response Length Analysis
Average response length distribution across Search Arena models is shown in Figure D.6 (Left). Reasoning models tend to be more verbose except for sonar-reasoning. sonar-pro’s version with a higher search context generates longer responses compared to the version with a medium context. Response length distribution is uniform across OpenAI’s models.
Figure D.6: (Left) Response length distribution. Reasoning models and models with high search context size tend to be more verbose. (Right) Citation count distribution. As expected, models with higher search context size cite more sources. Reasoning models cite fewer sources compared to non-reasoning variants (e.g., sonar-reasoning-pro-high vs sonar-pro-high).
We control for response length difference in the Bradley-Terry model and compute the corresponding coefficient across different subsets of the full data, broken down by intent category (see Figure D.7). For all intent categories, the coefficient (effect) is statistically significant; however, as expected, the effect is smallest for Factual Lookup prompts.
Figure D.7: Response Length Control across Intents. Bradley-Terry coefficients corresponding to response length across different intent categories. Length has less effect on Factual Lookup prompts.
@@ E.1 Citation Count
Citation count distribution is shown in Figure D.6 (Right). As expected, models with higher search context size cite more sources (e.g., sonar-pro-high vs sonar-pro). Furthermore, reasoning models tend to cite less sources compared to non-reasoning variants (e.g., sonar-reasoning-pro-high vs sonar-pro-high, sonar-reasoning vs sonar). We hypothesize that reasoning models synthesize and filter irrelevant sources before final response generation, resulting in less cited sources in the final response (see Figure 3). Interestingly, gemini-2.5-pro-grounding cites fewer sources compared to gemini-2.5-flash-grounding, suggesting that even though both are reasoning models, gemini-2.5-pro-grounding filters out more sources from the final response.
We then control for citation count difference in the Bradley-Terry model and compute the corresponding coefficient across different subsets of the dataset broken down by intent category (see Figure E.1). The effect of citation count on Guidance (e.g., debugging, problem-solving) prompts is not significant. The effect is largest on Analysis prompts.
Figure E.1: Citation Count Control across Intents. Bradley-Terry coefficients corresponding to citation count across different intent categories. Citation count does not have significant effect on Guidance (e.g., debugging, problem-solving) prompts. The effect is largest for Analysis queries.
@@ E.2 Citation Sources
Figure E.2: Citation Source Distribution across Model Families. Models from different providers are biased towards different types of sources: (1) Perplexity models prefer citing YouTube, social media, and community blogs, (2) OpenAI models are biased towards mainstream news outlets.
Figure E.3: Citation Source Distribution across Non-reasoning (sonar-pro-high) and Reasoning (sonar-reasoning-pro-high) Models. Although both systems are based on the same search module, the reasoning model (dark blue) cites more community and blog sources and less Wikipedia articles compared to the non-reasoning variant (light blue).
The distribution of citation domain categories is shown in Figure E.2. Perplexity’s models tend to cite social media (including YouTube) and community platforms (e.g., Reddit, Quora). OpenAI’s GPT-4o-based models are more biased towards mainstream news outlets. Google’s Gemini models are in between. To further understand the effect of reasoning, we analyze the distributional differences of the cited sources by sonar-pro-high (non-reasoning) and sonar-reasoning-pro-high (reasoning), as reported in Figure E.3. We find that even though both systems are based on the same search module, the reasoning model cites more community and blog sources and less Wikipedia articles, compared to the non-reasoning model. These results are consistent with our correlational analysis of user preferences and cited domains (e.g., users rejecting Wikipedia sources in specific settings).
We can also control for all features simultaneously (response length, number of citations, and citation sources) to compute adjusted model scores. The scores and rankings before and after applying these controls are shown in Figure E.4. We observe that the model scores and rankings tend to converge after the controls are applied. This convergence is particularly evident within model families: the top three models from Perplexity show significant convergence. This suggests that much of the variation between models in the same family is accounted for by the control features. We note that features such as factuality and coverage may be correlated with the number of citations or specific citation sources; we leave this analysis to future work.
Figure E.4: Model Scores Before and After Control. Model scores and rankings converge after the controls are applied.
@@ E.3 Citation Attribution Analysis
LLM Prompt for Claim–Citation Attribution Assessment Task: Tell me if the claim can be verified or supported by the web content. Specifically: If the claim is supported by the web content, return "support" If the claim is contradict to the web content, return "contradict" If the claim is completely irrelevant to the web content, return "irrelevant" Claim: {claim} Web content: {web content} Return in the json format: { "reasoning": "<explain the reasoning>", "answer": <"support" | "contradict" | "irrelevant"> }
@@ Appendix F Cross-Setting Analysis
For offline evaluation of models in search settings, we use BrowseComp (Wei et al., 2025) and SimpleQA (Wei et al., 2024) [^2] Due to the high cost of running search-augmented LLMs, we evaluated each model on the same randomly sampled subset of 500 questions. The subset was sampled once and shared across all models.. For testing the general performance of the models in non-search settings, we use ArenaHard-v2 (Li et al., 2024). BrowseComp and SimpleQA assess factuality based on well-specified ground truths, while ArenaHard-v2 utilizes an LLM-as-a-judge framework on a set of challenging prompts filtered from Chatbot Arena’s Text Arena dataset. Although these benchmarks test different aspects of standard and search-augmented LLMs, all are partially captured by the Search Arena prompt distribution shown in Figure 1.
@@
The selected models have near-zero accuracy on the BrowseComp benchmark; this finding is not surprising, as the questions are specifically designed to evaluate and challenge deep research pipelines. The model scores for the rest of the benchmarks–Search Arena, SimpleQA, and ArenaHard-v2–are provided in Table F.1. Additionally, we calculate model win rates on two subsets of the Search Arena dataset based on the annotated intent classes (see Section 2)–Search Arena (Fact+Synth) includes only Factual Lookup and Info Synthesis queries, while Search Arena (Other) contains the remaining subset. We use Kendall’s tau for comparing ranks and Pearson correlation for comparing raw scores across benchmarks. On SimpleQA, model accuracy is saturated (ranging from 89% to 93%), with minimal separability between the models and low agreement with Search Arena ($\tau=0.422$, $r=0.582$). In contrast, ArenaHard-v2 shows greater performance variance and separability, and has a higher agreement ($\tau=0.556$, $r=0.844$) with Search Arena compared to SimpleQA. However, the rankings differ; notably, the three reasoning models rank lower in ArenaHard-v2, suggesting that, while combining web search with reasoning improves performance in Search Arena, it may degrade performance on ArenaHard-v2 prompts. Additionally, when comparing the two subsets of the Search Arena with ArenaHard, Search Arena (Other) has higher agreement ($\tau=0.644$, $r=0.882$) compared to Search Arena (Fact+Synth) ($\tau=0.511$, $r=0.675$). This finding is expected as the prompt distribution in Search Arena (Other) is closer to that of ArenaHard (e.g., creative writing, problem-solving).
@@
Furthermore, we compare model performance with and without search on the SimpleQA and ArenaHard-v2 benchmarks. We used Gemini models in this case study, as search is implemented as a tool (GoogleSearch tool) and can be easily turned on and off. The performance change is shown in Figure F.2. Search significantly improves model performance on SimpleQA (models’ performance saturates at around 90%). However, performance degrades on ArenaHard-v2, with non-search variants achieving higher scores.
@@
Table F.1: Model scores across three benchmarks–Search Arena (win rate), SimpleQA (accuracy), and ArenaHard-v2 (win rate). On SimpleQA, models’ performance is saturated and lead to minimal separability in model scores. Search Arena and ArenaHard-v2 provide more separability, but the rankings are different.
@@
Figure F.1: Model Scores Across Benchmark. (1) Models’ performance on SimpleQA is saturated with minimal separability. (2) Performance variance in ArenaHard-v2 is comparable to that of Search Arena, but the rankings are different, with reasoning models having higher performance in search settings. (3) Model scores and ordering in Search Arena (Other) is closer to that of ArenaHard compared to Search Arena (Fact+Synth).
Figure F.2: SimpleQA vs ArenaHard-v2 Performance. Search and non-search performance of Gemini models on SimpleQA and ArenaHard-v2 benchmarks. Search improves performance on SimpleQA, while degrades the score on ArenaHard-v2.
Appendix G Additional Comparative Analyses
G.1 CORAL
As our setup is similar to prior work in conversational search, we conduct additional comparative analysis with CORAL (Cheng et al., 2025a), a benchmark designed to assess RAG systems in multi-turn conversational settings. Compared to CORAL, Search Arena conversations are collected from natural open-ended interactions between humans and search-augmented LLMs and contain user-provided preference labels.
Similar to the datasets studied in the main paper, CORAL focuses on English-only conversations, whereas the Search Arena prompts cover a diverse set of 71 languages. Search Arena (24k) also is larger in scale compared to CORAL (8k). Furthermore, in contrast to SimpleQA (Wei et al., 2024) and BrowseComp (Wei et al., 2025) and similar to Search Arena, CORAL includes multi-turn conversations. Compared to Search Arena (Figure B.3), CORAL conversations are longer on average (Figure G.1, Left). In terms of prompts, CORAL focuses on information-seeking queries only, whereas Search Arena covers a broader range of intents (e.g., factual lookup, creative generation). The prompt length distribution of Search Arena prompts is also more varied compared to CORAL, with the latter containing more concise questions (see Figure G.1, Right).
Additionally, in contrast to Search Arena, CORAL explicitly constructs ground truth answers, assuming a bounded set of documents (including golden). Search Arena focuses on human preferences, which, as observed in Section 3, do not need to correlate with factuality and quality of the retrieval. Finally, Search Arena data is collected from real-world natural interactions, enabling us to study real-world usage of the models and analyze how model response features correlate with the users’ preferences. Thus, the differences between the two datasets are due to the differences in the corresponding intents, use cases, and motivations.
G.2 WildChat
Similar to the Text Arena and Search Arena datasets studied in the main paper, WildChat (Zhao et al., 2024) is a crowd-sourced human-LLM conversation dataset. Similar to the Text Arena and in contrast to the Search Arena, the interactions in WildChat are with regular LLMs without access to additional tools, such as web search.
Despite the differences in the settings, we compare WildChat with Search Arena across multiple axes; we use a random sample (100k) of the non-toxic filtered version (3.2M) of the original dataset for the comparative analysis below. First, the full WildChat dataset is much larger in scale (4.8M), compared to Search Arena (24k). Both datasets are multilingual, containing interactions with over 70 languages. Turn and prompt length distributions of the WildChat dataset are provided in Figure G.2. WildChat turn count distribution is similar to that of Search Arena, with Search Arena containing slightly more multi-turn conversations. Additionally, WildChat prompts are longer on average (594.9) compared to those of Search Arena (57.1). This difference arises from the difference in the user intents across the two datasets: as reported in the Text Arena comparative analysis, search-enabled settings significantly impact the intent of the queries (e.g., users more likely to ask real-time information-seeking questions).
Figure G.1: (Left) CORAL conversation length (number of turns) distribution; CORAL conversations are longer than those in Search Arena, with 8.3 turns on average. (Right) CORAL prompt length distribution; CORAL prompts (red) are more concise and less diverse (lower variance) compared to Search Arena prompts (blue).
Figure G.2: (Left) WildChat conversation length (number of turns) distribution; WildChat conversations are similar in distribution to Search Arena Figure B.3, with slightly fewer multi-turn samples. (Right) WildChat prompt length distribution; WildChat prompts (pink) are longer compared to those of Search Arena (blue).